Constants Part A An airplane pilot wishes to fly due west. A wind of 84.0 km/h is blowing toward the south the airspood of the plane Ots speed in sti. air) is 400.0 km/h . in which direction should the pilot head? Express your answer as an angle measured north of west north of west Submit Request Answer Part B What is the speed of the plane over the ground? km/h Submit Request Answer Provide Feedback
Answer #1

the planes ground speed

V=\sqrt{} airspeed2-windspeed2=\sqrt{}4002-842=391km/h

and the angle , to head in the west direction the the north component of the velocity must match the speed of the wind to the south.

tan\Theta=84/391

\Theta=tan-1(84/391)=12o

so answer A) 12o or 12.1o or 12.12o b) 391 km/h or 391.1 km/h